# A Convex Hull Cheapest Insertion Heuristic for the Non-Euclidean TSP

### v3

@@ 5.2 Performance Analysis
The first column of Table 2 lists the name of the instance, each formatted with an alphabetic prefix followed by a numeric value that indicates the number of points in that instance. The cause of its non-Euclidean characteristic is listed alongside the instance name, while the second column provides the resulting DF. It can be seen that the DF is positively correlated with the number of separators, though it is also dependent on the spatial distribution of points and their relative position with the added separators. The remaining columns list the solutions found using the various approaches, with the mean ($\mu$) and standard deviation ($\sigma$) listed for the metaheuristic solutions found at the end of one minute.
@@
Figure 6: Performance analysis of the ACHCI algorithm and heuristics for TSP problems without precedence constraints.
@@
To visualize the effect of DF, the ratio of tour costs obtained using the NN and NI heuristic to that obtained using the ACHCI algorithm is shown in Fig. LABEL:H_ratio_DF for each of the 54 TSPLIB instances studied. It is clear that the ACHCI algorithm largely outperforms these competing heuristics, indicating that the proposed heuristic is well suited to a variety of non-Euclidean point cloud configurations. For increasing DF, a diminishing trend is observed in the advantage that the ACHCI algorithm provides, which is attributed to the increasing degree of approximation caused by the MDS when projecting points from an increasingly distorted non-Euclidean space to a Euclidean 2D space. It is also seen in Fig. LABEL:H_ratio_points that the relative performance of the ACHCI heuristic is largely unaffected by the number of points. The ACHCI tour cost was lower than that of the NN and NI heuristic solutions in 99% and 90% of the cases studied.
When comparing the performance of metaheuristic approaches, box plots enable the visualization of the stochasticity involved in the final solutions found. As expected, the relative performance of the GA and ACO solutions with respect to the ACHCI solutions decrease with increasing number of points as seen in Fig. LABEL:MH_ratio_points. This is because of the exponentially larger search space involved, and while the GA outputs solutions for any number of points because of the random generation of solutions, the ACO does not produce any solutions for larger instances. This is attributed to the resource intensive operation of constructing each tour by sampling each edge tour from the initial transition probabilities. It can be seen however that the performance of GA and ACO for smaller problem sizes is comparable and the relative performance is mostly unaffected by the DF, as seen in Fig. LABEL:MH_ratio_DF. When accounting for all problem sizes and DFs, the ACHCI cost outperforms the mean GA and ACO costs in 87 % and 95 % of the cases overall.
A worst case complexity of $O(n^{3})$ characterizes the ACHCI heuristic, as seen in Fig. LABEL:time where the x axis is the number of points raised to $3$. This is attributed to the eigenvalue decomposition involved with MDS and the cheapest insertion criteria used when selecting points while building the tour. Because of the greedy nature of the NN and NI heuristics, they almost instantaneously provides tours regardless of the number of points, and it is therefore worthwhile to compute the NN and NI tours in addition to the ACHCI tour, to simply choose the tour with minimum cost between the three.
@@ 6 Conclusion
The CHCI algorithm has been adapted for non-Euclidean instances of the TSP by leveraging MDS to approximate non-Euclidean points into Euclidean space so as to compute the initializing convex hull set of points. Empirical evaluations demonstrate that the ACHCI heuristic offers significant improvements over both heuristic and metaheuristic algorithms when applied to TSPs in non-Euclidean spaces.

### v5

@@ 5.2 Performance Analysis
The first column of Table 2 lists the name of the instance, each formatted with an alphabetic prefix followed by a numeric value that indicates the number of points in that instance. The cause of its non-Euclidean characteristic is listed alongside the instance name, along with the resulting DF. It can be seen that the DF is positively correlated with the number of separators, though it is also dependent on the spatial distribution of points and their relative position with the added separators. The remaining columns list the solutions found using the various approaches, with the mean ($\mu$) and standard deviation ($\sigma$) listed for the metaheuristic solutions found at the end of one minute.
@@
Figure 6: Computation time comparison
@@
To visualize the effect of DF, the ratio of tour costs obtained using the NN and NI heuristic to that obtained using the ACHCI algorithm is shown in Fig. LABEL:H_ratio_DF for each of the TSPLIB instances studied. It is clear that the ACHCI algorithm largely outperforms these competing heuristics, indicating that the proposed heuristic is well suited to a variety of non-Euclidean point cloud configurations. For increasing DF, a diminishing trend is observed in the advantage that the ACHCI algorithm provides, which is attributed to the increasing degree of approximation caused by the MDS when projecting points from an increasingly distorted non-Euclidean space to a Euclidean 2D space. It is also seen in Fig. LABEL:H_ratio_points that the relative performance of the ACHCI heuristic is largely unaffected by the number of points. The ACHCI tour cost was lower than NN and NI heuristic solutions in 88% and 99% of the cases studied respectively.
A worst case complexity of $O(n^{3})$ characterizes the ACHCI heuristic, as seen in Fig. 6 where the x axis is the number of points raised to $3$. This is attributed to the eigenvalue decomposition involved with MDS and the cheapest insertion criteria used when selecting points while building the tour. The NI heuristic also uses the cheapest insertion criteria but requires more computation time than the ACHCI heuristic because it is not initiated with the convex hull of points, resulting in an increased number of points for insertion. This results in no solution being found within the 60 second time limit for larger problem instances, as seen in Table 2. Because of the greedy nature of the NN heuristic, tours are produced almost instantaneously regardless of the number of points.
When comparing the performance of metaheuristic approaches, box plots enable the visualization of the stochasticity involved in the final solutions found. The relative performance of the GA and ACO solutions with respect to the ACHCI solutions decrease with increasing number of points as seen in Fig. LABEL:MH_ratio_points. This is because of the exponentially larger search space involved, and while the GA outputs solutions for any number of points because of the random generation of solutions, the ACO does not produce any solutions for larger instances. This is attributed to the resource intensive operation of constructing each ACO tour by sampling each edge from the initial transition probabilities. It can be seen however that the performance of GA and ACO for smaller problem sizes is comparable and the relative performance is mostly unaffected by the DF, as seen in Fig. LABEL:MH_ratio_DF. The ACHCI cost is lower than the mean GA and ACO costs in 87 % and 95 % of the studied test cases respectively.
@@ 6 Conclusion
The utilization of MDS enables the ACHCI heuristic to capitalize on the property that points on the boundary of the convex hull are visited in the same cyclic order as they appear in their optimal Euclidean TSP tour. While the initiating convex hull is based on an approximation of points in a Euclidean space, it is seen empirically that the non-Euclidean tours generated by the ACHCI heuristic outperform the benchmark algorithms. When compared with metaheuristic algorithms like the GA and ACO that iteratively update a population of solutions at high computational expense, the ACHCI tours outperformed them in a vast majority of cases, especially in larger test instances. ACHCI tour costs were lower than NI heuristic tour costs in 99% of the test cases studied, while also providing solutions more rapidly due to the reduced number of points for insertion. Though the NN heuristic is faster than the ACHCI heuristic because of its greedy approach to tour construction, the ACHCI tour costs were lower in 89% of the test cases. Considering the near instantaneous generation of NN tours, it may therefore be worthwhile to compute the NN tours in addition to the ACHCI tour, and simply choose the tour with the minimum cost.
Computational experiments presented in this paper were based on non-Euclidean test cases derived from TSPLIB instances by either the presence of obstacles, or the utilization of the $\mathcal{L}_{1}$ norm. While both these modifications are based on practical applications, it is easy to see that a similar performance can be expected from other real-world requirements that may arise out of other cost functions such as fuel consumption, emissions, travel time or driver comfort.